OpenStudy (anonymous):

Find the slope of a line parallel to y= -2/5x+4/5

OpenStudy (sleepyjess):

If it is parallel, that means the lines have the same slope.

OpenStudy (anonymous):

your slope is m Y=Mx + B

OpenStudy (adamk):

So the answer is the same "m" term in y = mx+b

OpenStudy (anonymous):

so what is M in your equation?

OpenStudy (nurali):

y=mx+b

OpenStudy (nurali):

m=slope

OpenStudy (anonymous):

@crystalm

OpenStudy (nurali):

y=-2/5x+4/5

OpenStudy (nurali):

m=?

OpenStudy (anonymous):

I don't get how to do this at all, this is why i needed help. :(

OpenStudy (anonymous):

Pm For more help @crystalm

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

hey you still there @crystalm